OpenStudy (anonymous):

80% of a number is 96. What is 25% of that same number? A. 20 B. 30 C. 40 D. 120

OpenStudy (anonymous):

Maybe or b I do believe

OpenStudy (elyssac):

96/80=x/25 96(25)=80x 2400=80x 2400/80=x=30 x=30

OpenStudy (anonymous):

Okay, so first you put 80/100 because 80 percent as a fraction is 80/100. then you put x/25 you put x because x is what you're trying to find out. Then you do 80/100 x/25 you do it across like 25 x 80 and 100 multiplied by x but you can't because you're trying to figure out what x is so you do 100 x 80 x 25 (you just do 80 x 25) Then, you do what ever 80 x 25 equals, divided by 100. So, 80 x 25 = 2,000. Now you do 2,000 divided by 100. Then you'll get your answer 2,000 divided by 100 is 20.

OpenStudy (anonymous):

IT's a.

OpenStudy (anonymous):

It's A.

OpenStudy (anonymous):

My bad

OpenStudy (anonymous):

It's alright! :)

OpenStudy (anonymous):

Medal?